in a cilender8 inches in diameter and12 inches in length. another cilender inside the cilinder with a clesrence of 1 inch all around all sides.and closed to seal the top compleatly. on the bottom a small hole to inkect liquid nitrogin.pour liuqid into container,inject liquid nitrogin. how many seconds will it take to freeze liquid? such as water etc.
1
Expert's answer
2013-06-26T07:39:57-0400
The volume of water will be calculated by formula:
V = 3.1416 * r^2 * h
so, V = 4 x 4 x 3.1416 x 12 = 603.1872 cubic inches or 9.88 L of water will take to freeze when it starts at 25°C.

So energy needed to freeze is:
E = cm∆T
m = V , because the density of water is 1.
m = 9.88 kg
E = 4.17 J g-1 K-1 x 9880g x 25°C = 1029990 J
The amount of energy, that needed to freeze the water in one second is 60 J
So if we have 1029990 J of energy, we can calculate the time that needed to freeze the 9,88 L of water
It will be 1029990/60 = 17166.5 seconds
Answer: 17166.5 seconds will it take to freeze water